OpenStudy (anonymous):

Three vertices of a rectangle are given. Find the coordinates of the fourth vertex. (0, 5), (5, 0), (2, –3) A. (–3, 2) B. (2, –3) C. (3, 2) D. (3, –2)

OpenStudy (anonymous):

well i give up good bye

OpenStudy (anonymous):

@SuperflyTyGuy

OpenStudy (superflytyguy):

A

OpenStudy (superflytyguy):

A. 100%

OpenStudy (anonymous):

k i have more

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

ok lets start working

OpenStudy (aaliyahfairgood):

k ill draw it out

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i have a diffrent one i need help with

OpenStudy (anonymous):

Three vertices of a rectangle are given. Find the coordinates of the fourth vertex. (–3, –3), (7, –3), (7, 2) A. (2, –3) B. (3, 2) C. (3, –2) D. (–3, 2)

OpenStudy (anonymous):

do you know the answer

OpenStudy (superflytyguy):

one sec

OpenStudy (superflytyguy):

D. -3,2

OpenStudy (anonymous):

k wait i have more

OpenStudy (aaliyahfairgood):

|0,5) (5,0) D=sq. root (5-0)+(0-0) D=sq. root (5)+(0) D=sq. root (5) (5,0) (2,-3) D=sq. root (2-5)+ (-3-0) D=sq. root (-3) + (-3) D=sq. root (-3-3) D=sq. root (-6)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so d

OpenStudy (anonymous):

@aaliyahfairgood

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

@aaliyahfairgood

OpenStudy (anonymous):

Three vertices of a square are (–1, 3), (–1, –1), and (3, 3). What is the fourth vertex of the square? A. (–1, –3) B. (1, 3) C. (3, 0) D. (3, –1)